satellite73 (satellite73):

find \(k\) so that \[2^k+2^{34}+2^{37}\] is a perfect square

OpenStudy (anonymous):

i have an answer, want to know if there is another method

OpenStudy (anonymous):

let k=34+y then, 2^(34+y) +2^34 +2^37 = 2^34 ( 2^y +9)

OpenStudy (anonymous):

I guess that would simplify the question

OpenStudy (anonymous):

so y= 4

OpenStudy (anonymous):

yeah that is pretty much the same method i used, thanks

OpenStudy (anonymous):

thus k = 34+4=38

OpenStudy (anonymous):

wrote \(2^{34}(1+2^3+2^n)\) and found \(n=4\) i was wondering if there was some kind of other reasoning

OpenStudy (jhannybean):

I finally understood this. lol (after being afk).